An expandable balloon contains 3.50 dm3 of helium at 101. kPa and 23.0oC. What volume will the balloon occupy when the pressure drops to 85.0 kPa and the temperature is 10.0oC?


1
Expert's answer
2021-06-24T11:32:02-0400

"\frac{P_1×V_1}{T_1}=\frac{P_2×V_2}{T_2}"


"\frac{101×3.5}{296}=\frac{85×V_2}{283}"


"V_2=3.98dm^3"